# Effects of dynamical friction on perturbations for evolving dark energy

### v1

@@ A.3 Corrections to the halo mass function
When dark energy clusters, also the halo mass function is modified. To take into accounts these modifications, we follow the work by Creminelli et al. (2010). According to the authors, one can scale the halo mass function calculated for a smooth model with the Sheth-Tormen Sheth et al. (2001) formulation with the halo mass function calculated with the Press-Schechter Press and Schechter (1974) formalism.
The halo mass function for smooth models is thus Creminelli et al. (2010)
@@
where it is understood that when the particular dark energy is clustering, one needs to use the appropriate matter power spectrum and linear overdensity parameter $\delta_{\rm c}$. The reason for this choice is that the ratio of the Press-Schechter functions will mitigate its shortcomings. We present our results in Figs. 14 and 15 for dynamical and dissipative pressure models, respectively. We consider the ration between the clustering model with the corresponding smooth model.
@@
For dynamical friction, we notice that the dominant contribution is that of clustering. Whether we predict more or less objects depends on the model considered. If the model has a quintessence equation of state, then clustering predicts more objects, with differences ranging from subpercent level for models 1 and 5 to a couple of percent for model 3 for masses of the order $10^{13}\,M_{\odot}\,h^{-1}$. At higher masses differences can be more than 10%. We notice that these substantial differences appear for model 3, whose equation of state is the most distant from that of the $\Lambda$CDM model. For phantom models, we observe a decrease in structure formation, general of a few percent and more substantial for model 6, as it is the model more different from $\Lambda$CDM.
@@
Very similar results, from both a qualitative and a quantitative point of view, are obtained for dissipative pressure friction. Also in this case the clustering is the dominating effect and difference are maximised for models 3 and 6. Except for the latter, models with different values of $\Gamma$ behave very similarly to each other.
@@
We also notice that our results are in good qualitative agreement with the original work of Creminelli et al. (2010).
Appendix B Derivation of the equation for the radius $R$
In this section we discuss how we can recover Eq. (8) from the fluid approach, but first we start from the equation of the radius $R$ and derive a second order differential equation for the matter overdensity $\delta$. We limit ourselves to the case of smooth dark energy, as that of clustering dark energy is totally analogous and will simply modify the Poisson equation with a term proportional to $\delta_{\rm de}$. To this purpose, we use the relation Engineer et al. (2000); Pace et al. (2019); Pace and Schimd (2022)
$1+\delta=\lambda\frac{a^{3}}{R^{3}}\,,$ (49)
with $\lambda$ constant and $a$ the scale factor.
Starting from Eq. (8), we need an expression for the velocity and the acceleration, which read, respectively,
$\frac{\dot{R}}{R}=$ $\,\frac{1}{3}\left[3H-\frac{\dot{\delta}}{1+\delta}\right]\,,$
$\frac{\ddot{R}}{R}=$ $\,\frac{1}{9}\left[3H-\frac{\dot{\delta}}{1+\delta}\right]^{2}+\frac{1}{3}\left[3\dot{H}+\frac{\dot{\delta}^{2}}{(1+\delta)^{2}}-\frac{\ddot{\delta}}{1+\delta}\right]\,.$
We will also use the Friedmann equation $H^{2}+\dot{H}=-\tfrac{4\pi G}{3}[\bar{\rho}_{\rm m}+\bar{\rho}_{\rm de}(1+3w_{\rm de})]$ with $G$ Newton’s gravitational constant.
With these equations, after a few algebraic steps one can easily find Eq. (II.1), which, as extensively discussed in the main body of this work, leads to problems.
We want, now, to derive the equation for $\delta$ using the fluid approach, by suitably modifying the Euler equation. This will allow us to highlight where the problems we discussed above originate.
Following Antonuccio-Delogu and Colafrancesco (1994), we use the following expression for the Euler equation [^10] Differently from Antonuccio-Delogu and Colafrancesco (1994), though, we include also nonlinear terms.
$\left.\frac{\partial\mathbf{v}}{\partial t}\right|_{\mathbf{r}}+\left(\mathbf{v}\cdot\nabla_{\mathbf{r}}\right)\mathbf{v}+\nabla_{\mathbf{r}}\Phi+\eta\mathbf{v}=0\,,$ (50)
which describes the macroscopic velocity of a fluid as in the Langevin approximation for the Brownian motion. In Eq. (50), derivatives are done with respect to the physical coordinate $\mathbf{r}$. The gravitation potential $\Phi$ consists of two terms: $\Phi_{0}$ related to a uniform matter distribution and such that $\nabla_{\mathbf{x}}\Phi_{0}=-a\ddot{a}\mathbf{x}$ and $\phi$ associated to inhomogeneities.
We now move from the time derivative with fixed physical coordinate $\mathbf{r}$ to time derivative with fixed comoving coordinate $\mathbf{x}$ using the well known relation
$\left.\frac{\partial}{\partial t}\right|_{\mathbf{r}}=\left.\frac{\partial}{\partial t}\right|_{\mathbf{x}}-H\mathbf{x}\cdot\nabla_{\mathbf{x}}\,,$
and write $\mathbf{r}=a\mathbf{x}$ and express the velocity $\mathbf{v}$ in terms of the comoving peculiar velocity $\mathbf{u}$: $\mathbf{v}=aH\mathbf{x}+a\mathbf{u}$.
Inserting these expressions, the Euler equation for the peculiar velocity is now
$\left.\frac{\partial\mathbf{u}}{\partial t}\right|_{\mathbf{x}}+(2H+\eta)\mathbf{u}+\eta H\mathbf{x}+\left(\mathbf{u}\cdot\nabla_{\mathbf{x}}\right)\mathbf{u}+\frac{1}{a^{2}}\nabla_{\mathbf{x}}\phi=0\,.$ (51)
Note how this equation contains an additional term, $\eta H\mathbf{x}$, which only depends on background quantities. It is exactly this term the source of the unphysical behaviour of the growth factor we discussed above. It is now clear to see how this term originates from using physical coordinates rather than comoving coordinates. Also, the above derivation implies that dynamical friction affects equally background and perturbations, which is though not the case.
From now on, we omit, for brevity, the subscript $\mathbf{x}$ and we imply that derivatives are done with respect to the comoving coordinate.
Taking the divergence of Eq. (51) we arrive to
$\dot{\theta}+\left(2H+\eta\right)\theta+3H\eta+\frac{1}{3}\theta^{2}+\sigma^{2}-\omega^{2}+\frac{1}{a^{2}}\nabla^{2}\phi=0\,,$ (52)
where $\theta=\nabla_{\mathbf{x}}\cdot\mathbf{u}$ is the divergence of the peculiar velocity, $\sigma$ the shear and $\omega$ the rotation.
The last step in our derivation is to use $\ln{a}$ as the independent variable and to write the Euler equation for the dimensionless variable $\tilde{\theta}=\theta/H$:
$\tilde{\theta}^{\prime}+\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\tilde{\theta}+\frac{1}{3}\tilde{\theta}^{2}+\tilde{\sigma}^{2}-\tilde{\omega}^{2}+3\frac{\eta}{H}+\frac{\nabla^{2}\phi}{a^{2}H^{2}}=0\,.$ (53)
It is easy to see that combining the continuity, Euler and Poisson equations we recover Eq. (II.1).
Dynamical friction thus enters in the equation of motion for perturbations as an additional source of friction in the Euler equation, with the result of modifying the coefficient of $\delta^{\prime}$ and adding an additional term to the overall system. This implies that also the linearised version of Eq. (II.1), which represents the growth factor, is modified and friction, therefore, also affects linear dynamics. The modified growth factor equation reads
$\delta^{\prime\prime}+\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\delta^{\prime}-\frac{3}{2}\Omega_{\rm m}(t)\delta-\frac{3\eta}{H}=0\,.$ (54)

### v2

@@ A.3 Corrections to the halo mass function
When dark energy clusters, the halo mass function is also modified. To take into account these modifications, we follow the work of [100]. According to the authors, one can scale the halo mass function calculated for a smooth model with the Sheth-Tormen [99] formulation with the halo mass function calculated with the Press-Schechter [98] formalism.
The halo mass function for smooth models is thus [100]
@@
where it is understood that when the particular dark energy model is clustering, one needs to use the appropriate matter power spectrum and linear overdensity parameter $\delta_{\rm c}$. The reason for this choice is that the ratio of the Press-Schechter functions will mitigate its shortcomings. We present our results in Fig. 11, where we consider the ratio between the clustering model and the corresponding smooth model.
@@
For dynamical friction, we notice that the dominant contribution is that of clustering. Whether we predict more or less objects depends on the model considered. If the model has a quintessence equation of state, then clustering predicts more objects, with differences ranging from the subpercent level for models DE1 and DE5 to a couple of percent for model DE3 for masses of the order $10^{13}\,M_{\odot}\,h^{-1}$. At higher masses, the differences can be greater than 10%. We notice that these substantial differences appear for model DE3, whose equation of state is the most distant from that of the $\Lambda$CDM model. For phantom models, we observe a decrease in structure formation, generally of a few percent and more substantial for model DE6, as it is the model more different from $\Lambda$CDM.
@@
Very similar results are obtained, from both a qualitative and a quantitative point of view, for dissipative pressure friction. Also in this case, clustering is the dominating effect and differences are maximized for models DE3 and DE6. Except for the latter, models with different values of $\Gamma$ behave very similarly to each other.
@@
We also notice that our results are qualitatively in good agreement with the original work of [100].
Figure 11: Number of objects above a given mass threshold $M$ for the model DE1, when dark energy is perturbed. We considered dynamical friction. Left panel shows results for dynamical friction, while the right panel for dissipative pressure friction. Line styles and colours are as in Fig. 1.
Data availability
No data was used for the research described in the article.